Explain whether it is possible for the amount of employed people to decrease while the unemployment rate decreases


1
Expert's answer
2021-05-11T14:21:08-0400

"Unemployment \space Rate = \frac{Total\space Unemployed}{Labor Force} \times100"


"Labor \space Force = Employed + Unemployed"


Suppose initially there were 25 employed and 50 unemployed people in a hypothetical economy.

"Labor\space Force = 25 + 50 = 75"

"Unemployment\space Rate = \frac{50}{75} \times 100 = 66.7\%"


Now, Suppose out of 25 employed 10 people opted out of labor force and due to some family commitments they stopped working and stopped actively looking for work. Then new employed figure will be,

25 - 10 = 15


Also suppose that out of 50 unemployed, 25 became discouraged and stopped looking for work actively. So , these will be out of labor force. New Unemployed figure = 50 - 25 = 25

"New\space Labor \space Force = 15 + 25 = 40"

"New \space Unemployment\space Rate = \frac{25}{40 }\times 100 = 62.5\%"


Thus, the number of employed has reduced and unemployment rate has also decreased. 

This means that when there is significant decline in labor force or working age population, then both number of employed and unemployment rate will decrease. Yes, it is possible for the amount of employed people to decrease while the unemployment rate decreases.